OpenStudy (anonymous):

find the limit of (-3x^3+x+1)/(2x^2-x+2) as x approaches -infinity

myininaya (myininaya):

divide top and bottom by x^2

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}\frac{\frac{-3x^3}{x^2}+\frac{x}{x^2}+\frac{1}{x^2}}{\frac{2x^2}{x^2}-\frac{x}{x^2}+\frac{2}{x^2}}\]

OpenStudy (anonymous):

oh myinaya thank you very much. i have alot of unique limit problems, would you be willing to help me?

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}\frac{-3x+\frac{1}{x}+\frac{1}{x^2}}{2-\frac{1}{x}+\frac{2}{x^2}}\] \[\frac{-3(x-> -\infty) +0+0}{2-0+0}=\infty\]